OpenStudy (anonymous):

7x^3/3 divided by 14x^2/6

OpenStudy (anonymous):

rewrite your question using the equation button below

OpenStudy (anonymous):

| | v

OpenStudy (anonymous):

\[7x ^{3}/3 \div 14x ^{2}/6\]

OpenStudy (anonymous):

okay. change the division sign to a multiplication sign by inverting the denominator and post what you get.

OpenStudy (anonymous):

\[\frac{\frac{7x^3}{3}}{\frac{14x^2}{6}}\]?

OpenStudy (anonymous):

\[\frac{7x^2}{3}*\frac{6}{14x^2}=\frac{42x^2}{42x^2}\]\[\frac{}{}\]=1

OpenStudy (anonymous):

\[\frac{a}{b}\div\frac{c}{d}=\frac{a}{b}*\frac{d}{c}\] so you have \[\frac{7x^3}{3}\div\frac{14x^2}{6}=\frac{7x^3}{3}*\frac{6}{14x^2}= ...\] just multiple the top lines together and the the bottoms, and simplify.

OpenStudy (anonymous):

I tried 1 and it doesnt work